OpenStudy (anonymous):

what is the standard form of the equation of the circle x^2-2x+y^2-8x+1=0

OpenStudy (jhannybean):

\[x^2 -2x+y^2-8x+1=0\] compelte the square for both x and y

OpenStudy (zzr0ck3r):

err 8y?

OpenStudy (zzr0ck3r):

its a race to see who can complete the square the fastest

OpenStudy (jhannybean):

\[x^2\color{red}{-2x}+y^2\color{red}{-8x}+1=0\]Combine like terms first.

OpenStudy (anonymous):

Ahahaha.

OpenStudy (jhannybean):

What is (-2-8)x?

OpenStudy (anonymous):

-10x right?

OpenStudy (jhannybean):

Yes.

OpenStudy (jhannybean):

\[x^2\color{red}{-2x}+y^2\color{red}{-8x}+1=0\]\[(x^2 \color{red}{-10x})+y^2 +1=0\]Now you want to complete the square inside the parenthesis by finding your new \(c\) value.

OpenStudy (jhannybean):

Basically you want to complete a quadratic inside the (), \(ax^2+bx+c\)

OpenStudy (jhannybean):

so finding \(c\) is easy, \(c=\left(\frac{b}{2}\right)^2 = \left(\frac{-10}{2}\right)^2 = (-5)^2 = 25\)

OpenStudy (jhannybean):

\[\color{blue}{c}=\left(\frac{b}{2}\right)^2 = \left(\frac{-10}{2}\right)^2 = (-5)^2 = \color{blue}{25}\]Now add that in .\[(x^2 \color{red}{-10x}+\color{blue}{25} )+y^2 +1-25=0\] Remember that when you add anything in to complete a square, you must also subtract by the same number outside the square.

OpenStudy (jhannybean):

Simplify it. \[(x-5)^2+y^2 -24=0\]Put the 24 onto the other side. \[(x-5)^2 +y^2 =-24\]

OpenStudy (jhannybean):

Standard form of a circle is: \((x-h)^2 +(y-k)^2 =r^2\)

OpenStudy (zzr0ck3r):

Are we sure that should not have been 8y not 8x?

OpenStudy (zzr0ck3r):

seems weird to test them on combining like variables at this stage...

OpenStudy (jhannybean):

We could do that too,I was just working with what was given, lol

OpenStudy (zzr0ck3r):

I know. ;) was more for the user to make sure, then you can see another way!

OpenStudy (anonymous):

wouldn't it be positive 24 not negative?

OpenStudy (jhannybean):

Oh yes you are right, \(1-25 = 0 \implies -24\) and on the other side of the 0 it becomes \(+24\)

OpenStudy (jhannybean):

Sorry.

OpenStudy (anonymous):

Naw your good it was very helpful!!! thank you!

OpenStudy (jhannybean):

i gotta go, when I come back I'll show you how to work the equation the way @zzr0ck3r mentioned. Or he can show you as well!

ganeshie8 (ganeshie8):

yeah i still think that -8x is a typo

OpenStudy (anonymous):

It's not... that is what is write on my paper.

OpenStudy (jhannybean):

To avoid simplifying your functon restricted only to x variables, your equation should have looked like: \( x^2-2x+y^2-8y+1=0\) You would have completed the square then for BOTH x and y.\[(x^2-2x)+(y^2-8y)=0\]

OpenStudy (jhannybean):

You should definitely try it this way, it'll give you practice in finding your center, as well as competing squares for 2 variables instead of just 1.

OpenStudy (jhannybean):

@zzr0ck3r it's your turn. Take care you guys!

OpenStudy (zzr0ck3r):

lol, if it was not 8y it don't matter....